OpenStudy (anonymous):

how can you solve, log base 6 of (x-3) - log base 6 of 4= log base 6 of (x+2)

OpenStudy (dumbcow):

use log property log x - logy = log(x/y) Then notice if logx = logy, then x=y --> (x-3)/4 = x+2

OpenStudy (anonymous):

how can i solve for X

OpenStudy (dumbcow):

its a linear equation, get the x's to one side multiply by 4 -> x-3 = 4(x+2)

OpenStudy (anonymous):

x would equal -11/3 right?

OpenStudy (anonymous):

thanks for the help